# Interplay of entanglement structures and stabilizer entropy in spin models

### v2

@@ 4.4 Cluster Ising model
A Symmetry-Protected Topological (SPT) phase is one that, as long as certain symmetries are present, is not adiabatically connected to a trivial product state. SPTs cannot be understood in the framework of local order parameters and spontaneous symmetry breaking. Instead, they are distinguished by nonlocal string-order parameters.
We consider infinite one-dimensional (1D) spin-$1/2$ chains described by the three-parameter Hamiltonian
$\hat{H}=\sum_{j}\left\{-g_{zz}\sigma^{z}_{j}\sigma^{z}_{j+1}-g_{x}\sigma^{x}_{j}+g_{zxz}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}\right\}$ (50)
This Hamiltonian is symmetric under global spin flips generated by $\prod_{j}\sigma^{x}_{j}$ as well as time-reversal (complex conjugation). Due to these symmetries, the model has a $Z_{2}\times Z_{2}$ SPT phase, as well as a trivial and a symmetry-broken phase.
Along the one-dimensional path parametrized as $g_{zz}=2(1-g^{2})$, $g_{x}=(1+g)^{2}$ and $g_{zxz}=(g-1)^{2}$ with tuning parameter $g$, the ground state of the model is described by exact MPS with bond dimension $\chi=2$ with the following tensors:
@@
This state was dubbed the MPS skeleton [136]. The MPS skeleton describes exactly the ground state even at the critical point $g=0$, despite the fact that the entanglement entropy there is strictly bounded by $\log(2)$. This is because the critical point is tricritical and there is no divergence of entanglement entropy with subsystem size, unlike in second-order Ising transitions. Since the ground state along the g-trajectory is a $\chi=2$ MPS, it is straightforward to calculate the SRE of the MPS ansatz [137]
@@
The transition between the trivial and the SPT phase occurs at the tricritical point between the three phases at $g=0$. For $g=\pm 1$ the SRE vanishes, $m_{2}=0$, implying the ground state at these points is a stabilizer state. For $g=-1$, we are deep in the SPT phase, and the ground state of the skeleton is the cluster state, which we know can be represented as a graph state and hence is a stabilizer state. For $g=+1$, we are deep in the trivial phase, hence its product ground state is also a stabilizer state. Finally, for $g=0$ we are located at the tricritical point where the ground state in the thermodynamic limit is a Greenberger-Horne-Zeilinger (GHZ) state [138], which is also a stabilizer state since it can be prepared using only Clifford gates. The more surprising feature is the peak of nonstabilizerness, which can be seen in Fig.11 (a). From Eq. (52). it is easy to show that the peak occurs at $g=\pm(3-2\sqrt{(}2))$, with the peak $m_{2}=0.3$, which is comparable to the SRE of the Ising model at its critical point.
In Fig. 11, we depict the behavior of $m_{2}$, the log-ratio $\log\Lambda$, and the capacity of entanglement $C_{E}$ as functions of the tuning parameter $g$ along a one-dimensional cut in the phase diagram of the cluster-Ising model. Data is shown for different system sizes to highlight size dependence. The first panel illustrates the behavior of the stabilizer Rényi entropy, confirming that $m_{2}$ exhibits two peaks at $g=\pm(3-2\sqrt{(}2))$, as predicted analytically, while vanishing at $g=\pm 1$ and at the tricritical point $g=0$, where the ground state is a stabilizer state. The second panel shows the antiflatness, which also peaks near the same values of $g$, indicating that the departure from stabilizerness is correlated with an increased complexity in the ground state wavefunction. Finally, the third panel presents the entanglement capacity $C_{E}$, which displays a strong peak around the phase transition, further reinforcing that this region is where quantum correlations are maximized.
Notably, all three quantities exhibit sharper features as the system size increases, suggesting that the peaks in nonstabilizerness and entanglement capacity behave similarly to critical singularities in a phase transition. These observations further highlight the capacity of magic as a diagnostic tool for characterizing quantum phase transitions, particularly in SPT systems where conventional order parameters fail.

### v3

@@ 4.4 Cluster Ising model
A Symmetry-Protected Topological (SPT) phase is characterized by the presence of a protecting symmetry that prevents the adiabatic deformation of the system’s ground state into a trivial product state. Unlike conventional phases, SPT phases cannot be described using local order parameters and are instead diagnosed using nonlocal observables, such as string order parameters.
We study a class of 1D spin-$1/2$ chain models with the following Hamiltonian:
$\hat{H}=\sum_{j}\left\{-g_{zz}\sigma^{z}_{j}\sigma^{z}_{j+1}-g_{x}\sigma^{x}_{j}+g_{zxz}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}\right\}.$ (50)
This model exhibits both global spin-flip symmetry (generated $\prod_{j}\sigma^{x}_{j}$) and time-reversal symmetry, which together protect a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT phase. The system also supports a trivial phase and a symmetry-broken phase.
We focus on a solvable trajectory in parameter space defined by the relations $g_{zz}=2(1-g^{2})$, $g_{x}=(1+g)^{2}$, and $g_{zxz}=(g-1)^{2}$. Along this path, the ground state admits an exact MPS description with bond dimension two [136]. The corresponding tensors are given by
@@
This variational state captures the ground state physics across all three phases, including the tricritical point at $g=0$, where three different phases meet. Notably, although the system is critical at $g=0$, the entanglement entropy remains finite and bounded by $\log 2$, reflecting the absence of conformal symmetry.
The SRE associated with this MPS can be computed exactly using transfer matrix techniques for translation-invariant systems. As shown in [137], the resulting expression is:
@@
This function displays a pronounced nonmonotonic behavior as $g$ varies, with local maxima at $g=\pm(3-2\sqrt{2})$, where the stabilizer entropy attains its peak value of approximately $0.3$. These peaks reflect enhanced non-Clifford correlations in the ground state. In contrast, the SRE vanishes at $g=\pm 1$ and $g=0$. For $g=1$, the ground state is a trivial product state, while for $g=-1$, it corresponds to a cluster state, a well-known stabilizer state within the SPT phase. At $g=0$, the system sits at the tricritical point and the thermodynamic ground state is a GHZ state [138], also a Clifford state with vanishing SRE.
Fig. 11 illustrates the dependence of $m_{2}$, the antiflatness measure $\log\Lambda$, and the entanglement capacity $C_{E}$ on the tuning parameter $g$. Data is shown for different system sizes to highlight size dependence. The first panel shows the behavior of the SRE. It presents a clear double-peak structure centered near $g=\pm(3-2\sqrt{2})$, with dips at the three stabilizer points discussed above. The second panel shows the antiflatness, which also peaks at $g=0$ and $g=(3-2\sqrt{2})$, indicating that the departure from stabilizerness is correlated with an increased complexity in the ground state wavefunction. Finally, the third panel presents the entanglement capacity $C_{E}$, which displays a strong peak around the phase transition, further reinforcing that this region is where quantum correlations are maximized.
Notably, all three quantities sharpen with increasing system size, indicating the emergence of critical-like scaling near the transition. This underlines the utility of nonstabilizerness as a probe of quantum phase transitions, particularly in scenarios where standard symmetry-breaking order parameters are absent.